# MMTU: A Massive Multi-Task Table Understanding and Reasoning Benchmark

### v2

@@ A.10 Knowledge-base mapping (KB Mapping)
Task Name Dataset Chat Model Reasoning Model
GPT-4o Llama33-70B o4-mini Deepseek-R1
Data-transform-reshape Auto-Tables 0.298 0.227 0.681 0.332
Transform-by-output-target-schema commercial-pipelines 0.077 0.154 0.231 0.077
github-pipelines 0.204 0.176 0.250 0.220
Average 0.140 0.165 0.241 0.148
Transform-by-input-output-table AutoPandas 0.593 0.519 0.815 0.815
Scythe 0.550 0.533 0.933 0.917
Average 0.571 0.526 0.874 0.866
Entity-Matching Amazon-Google 0.912 0.907 0.918 0.916
BeerAdvo-RateBeer 0.943 0.932 0.989 0.955
DBLP-ACM 0.992 0.975 0.988 0.990
DBLP-Scholar 0.962 0.961 0.968 0.957
Fodors-Zagats 0.989 0.989 0.995 0.989
Walmart-Amazon 0.967 0.949 0.968 0.974
iTunes-Amazon 0.915 0.953 0.991 0.887
Average 0.954 0.952 0.974 0.953
header-value-matching TableGPT 0.644 0.545 0.774 0.708
Schema-Matching DeepMDatasets 1.000 1.000 1.000 1.000
HXD 0.930 0.919 0.950 0.943
Wikidata 0.912 0.881 0.912 0.912
assays 0.825 0.818 0.902 0.924
miller2 0.924 0.947 0.936 0.916
prospect 0.919 0.893 0.986 0.972
Average 0.918 0.910 0.948 0.944
Data-Imputation WebTable 0.455 0.431 0.539 0.546
tablib 0.532 0.422 0.757 0.616
Average 0.494 0.426 0.648 0.581
Error-Detect Relational-Tables 0.266 0.176 0.168 0.140
Spreadsheet-Tables 0.188 0.152 0.120 0.078
Average 0.227 0.164 0.144 0.109
List-to-table TEGRA 0.612 0.552 0.584 0.531
semantic-join DataXFormer 0.676 0.640 0.819 0.730
SEMA-join 0.812 0.795 0.899 0.901
Average 0.744 0.718 0.859 0.815
equi-join-detect Auto-BI 0.567 0.576 0.655 0.691
Data-transform-pbe TDE 0.398 0.318 0.636 0.581
Transformation-text 0.467 0.349 0.651 0.630
Average 0.433 0.334 0.643 0.605
Formula-prediction-context cisco-random 0.065 0.032 0.175 0.127
enron-random 0.034 0.017 0.122 0.101
pge-random 0.030 0.050 0.113 0.130
ti-random 0.041 0.013 0.154 0.129
Average 0.042 0.028 0.141 0.122
semantic-transform DataXFormer 0.393 0.360 0.427 0.444
SEMA-join 0.861 0.865 0.877 0.863
Average 0.627 0.613 0.652 0.653
Arithmetic-Relationship Auto-Relate 0.221 0.284 0.702 0.638
Functional-Dependency Auto-Relate 0.683 0.702 0.689 0.741
String-Relationship Auto-Relate 0.021 0.007 0.405 0.485
Table-needle-in-a-haystack experiment8 0.441 0.154 0.884 0.650
Table-Locate-by-Row-Col experiment5 0.498 0.167 0.944 0.823
NL2SQL Archer 0.212 0.154 0.442 0.375
KaggleDBQA 0.292 0.492 0.508 0.503
Spider 0.732 0.766 0.774 0.756
WikiSQL 0.745 0.663 0.734 0.720
bird 0.405 0.393 0.478 0.456
Average 0.477 0.494 0.587 0.562
Table-QA FinQA 0.196 0.147 0.488 0.361
TableBench 0.410 0.420 0.575 0.554
WikiQA 0.682 0.653 0.819 0.769
Average 0.429 0.407 0.627 0.561
Table-Fact-Verification TabFact 0.841 0.794 0.943 0.922
Column-type-annotation SemTab2019 0.463 0.347 0.367 0.376
Columns-property-anotation SemTab2019 0.242 0.249 0.315 0.351
Cell-entity-annotation SemTab2019 0.673 0.610 0.705 0.738
Appendix B Detailed model performance at the dataset level
@@
Appendix C Long-context table understanding: A case study of “needle in a haystack in tables”
Figure 11: The needle-in-a-haystack test used in NLP context. Each cell in the heatmap corresponds to a test to retrieve a simple fact planted in a document, for a given document length and retrieval depth. Frontier models like GPT-4o and Gemini-2.5 can now typically achieve perfect accuracy, as shown by the all-green heatmap (results from Roberts et al. (2024)).
@@ Appendix C Long-context table understanding: A case study of “needle in a haystack in tables”
Recall that “Needle-in-a-haystack (NIH)” NIA ; Roberts et al. (2024), is a popular test traditionally used to evaluate long-context LLM’s ability to retrieve a simple fact (a needle) from a long document context (e.g., 100K or 1M tokens). Recent advances in frontier long-context LLMs have made this task almost irrelevant, as most frontier models such as GPT-4o, Llama-4, and Gemini-2.5 can now score perfectly on NIH Liu et al. (2024); Team et al. (2024); lla ; cla , like the heatmap in Figure 12 would show. Here the all-green heatmap indicates that this particular LLM under test, GPT-4o, can perfectly retrieve a needle planted at varying depth (y-axis), within documents of varying lengths (x-axis), with 100% accuracy.
@@
The results of NIHT are illustrated also using a heatmap in Figure 12. As we can see, compared to Figure 12, today’s frontier LLMs still face substantial challenges in correctly identifying a cell value within a two-dimensional table, like indicated by the red cells in the heatmap. This is especially true when we increase the number of columns (e.g., with more than 25 columns, LLM accuracy quickly falls below 0.5). Furthermore, we observe a strong asymmetry in the heatmap – e.g., with 50 rows and 5 columns we still have a reasonable accuracy of 0.81 (at the lower-left corner of the heatmap), but with 5 rows and 50 columns the accuracy drops to a lowly 0.43 (at the top-right corner), underscoring the challenge LLMs face when reading “vertically” in the column direction, which are consistent with our intuition that LLMs pre-trained on one-dimensional natural-language texts are less effective in reading two-dimensional tables “vertically”, like was also observed in Li et al. (2023c); Sui et al. (2024).

### v3

@@ Appendix C Additional Statistics
Appendix D Detailed model performance at the dataset level
@@
Appendix E Long-context table understanding: A case study of “needle in a haystack in tables”
Figure 10: The needle-in-a-haystack test used in NLP context. Each cell in the heatmap corresponds to a test to retrieve a simple fact planted in a document, for a given document length and retrieval depth. Frontier models like GPT-4o and Gemini-2.5 can now typically achieve perfect accuracy, as shown by the all-green heatmap (results from [niah-2]).
@@ Appendix E Long-context table understanding: A case study of “needle in a haystack in tables”
Recall that “Needle-in-a-haystack (NIH)” [NIAH-orig, niah-2], is a popular test traditionally used to evaluate long-context LLM’s ability to retrieve a simple fact (a needle) from a long document context (e.g., 100K or 1M tokens). Recent advances in frontier long-context LLMs have made this task almost irrelevant, as most frontier models such as GPT-4o, Llama-4, and Gemini-2.5 can now score perfectly on NIH [llm-deepseek-v3, gemini-niah, llama-4-release, claude-3-release], like the heatmap in Figure 11 would show. Here the all-green heatmap indicates that this particular LLM under test, GPT-4o, can perfectly retrieve a needle planted at varying depth (y-axis), within documents of varying lengths (x-axis), with 100% accuracy.
@@
The results of NIHT are illustrated also using a heatmap in Figure 11. As we can see, compared to Figure 11, today’s frontier LLMs still face substantial challenges in correctly identifying a cell value within a two-dimensional table, like indicated by the red cells in the heatmap. This is especially true when we increase the number of columns (e.g., with more than 25 columns, LLM accuracy quickly falls below 0.5). Furthermore, we observe a strong asymmetry in the heatmap – e.g., with 50 rows and 5 columns we still have a reasonable accuracy of 0.81 (at the lower-left corner of the heatmap), but with 5 rows and 50 columns the accuracy drops to a lowly 0.43 (at the top-right corner), underscoring the challenge LLMs face when reading “vertically” in the column direction, which are consistent with our intuition that LLMs pre-trained on one-dimensional natural-language texts are less effective in reading two-dimensional tables “vertically”, like was also observed in [table-gpt, table-llm].
Super-Categories Percentage
Reasoning & Coding Errors 40.3%
Table-Understanding Errors 25.6%
Knowledge Errors 16.4%
Other Errors 11.7%
Table 6: Distribution of Super-category Errors On Incorrect Questions by o4-mini
Super Categories Sub Categories Percentage
Table-Understanding Errors Row/Column index misalignment 0.283122
Reasoning & Coding Errors Incorrect reasoning 0.223264
Reasoning & Coding Errors Semantically incorrect code 0.138167
Knowledge Errors Incorrect knowledge 0.105523
Knowledge Errors Hallucinated facts 0.0334465
Other Errors Format error / Result-extraction failure 0.0323357
Other Errors Ground-truth ambiguity / quality 0.0323357
Table-Understanding Errors Complex table 0.0156742
Table-Understanding Errors Multi-table understanding 0.0153039
Knowledge Errors Numerical Errors 0.014008
Table 7: Distribution of Top-10 Sub-category Errors On Incorrect Questions by o4-mini.
Appendix F LLM-based Error analysis
Setup. We conducted a LLM-based error analysis to understand the underlying reasons of these errors on MMTU. We built a detailed categorization of typical errors that we observe in models’ answers, using the following 4 super-categories, and 14 sub-categories:
  • Super-category: Table-Understanding Errors
    • Row/Column index misalignment (e.g., the row/column index is not correctly recognized in a table, leading to incorrect model output)
    • Long context (e.g., long table causes model to produce incorrect results)
    • Multi-table understanding (e.g., join relationships between multiple tables are identified incorrectly, leading to incorrect answers)
    • Complex table (e.g., merged cells spanning multi-row/cols, hierarchical headers, etc., causing models to produce incorrect answers)
  • Super-category: Reasoning & Coding Errors
    • Syntax or execution error (if the task asks for SQL/Python code, and the code should encounter an execution error)
    • Semantically incorrect code (if the task asks for SQL/Python code, and the code is executable but the result is incorrect)
    • Incorrect reasoning (pick this category only if the model’s reasoning shows obvious flaws that directly lead to incorrect answers)
  • Super-category: Knowledge Errors
    • Hallucinated facts (e.g., in CEA, output knowledge base references that do not exist)
    • Incorrect knowledge (e.g., in data imputation, filled in a related but incorrect value)
    • Numerical Errors (e.g., in data imputation, filled a numerical value that is close but not precise)
  • Super-category: Other Errors
    • Format error / Result-extraction failure
    • Timeout
    • Ground-truth ambiguity / quality
    • None of the above: create a new category, and give a short description
Result. We used o4-mini to categorize all 11K questions for which o4-mini were incorrect on MMTU into these categories. The distribution of super-categories is shown in Table 6, note that the percentage does not sum to one due to format issue and LLM curated super-categories. The distribution of sub-categories is shown in Table 7. We observe that reasoning & coding, and table understanding remain the two largest categories of errors.